# Zero-shot sampling of adversarial entities in biomedical question answering

### v1

@@ 3 Datasets
To relate to real-world scenarios, we sourced the vocabulary datasets of drug and disease names from existing public databases. The drug names dataset (FDA-drugs) contains over 2.3k entities from known drugs approved by the United States Food and Drug Administration (FDA) and curated by Drug Central [^4] https://drugcentral.org/download (Ursu et al., 2017). FDA is a world authority that approves the legal distribution of drugs, therefore the dataset represents the drug-entity names encountered in daily life. The disease names dataset [^5] https://ctdbase.org/downloads/ (CTD-diseases) contains over 9.8k entities from the Comparative Toxicogenomics Database (Davis et al., 2009), which contains a comprehensive collection of uniquely documented chemical-gene-disease interactions for humans. Both databases are curated by domain experts through regular updates. The 2023 version of the datasets was used after minor data processing.
@@
We selected over 9.3k questions from the MedQA-USMLE (Jin et al., 2021) dataset and over 3.8k questions from the MedMCQA (Pal et al., 2022) dataset for benchmarking. Both datasets are entity-rich and cover a wide range of biomedical topics. The MedQA-USMLE dataset contains long-context questions that are used to assess medical students in the United States. The MedMCQA dataset contains short questions used in medical exams in India. They have been used in recent works for assessing the biomedical knowledge in LLMs (Singhal et al., 2023a; Han et al., 2023; Singhal et al., 2023b). Both datasets are publicly available and don’t contain personal information. We annotated the NEs in the two datasets according to the entity types of the Unified Medical Language System (UMLS) (Bodenreider, 2004) using scispaCy (Neumann et al., 2019). We then divided the QA datasets into drug- or disease-mention questions by the binary mention labels generated according to the UMLS entity types. (see Appendix C).
4 Attack implementation
4.1 Type-constrained entity selection
Implementing the framework in Section 2 and Appendix A for drug and disease names requires accounting for the entity types they are associated with. We used the entity type annotations according to the UMLS (Bodenreider, 2004) of the semantic groups Chemicals & Drugs and Disorders. For each question the model answers correctly, we selected the entity of the same semantic group in the distractors closest to the key as the one to attack. The choice of the closest distractor is to maintain consistency and is not uniquely associated with the effects observed here. Then, we selected the replacement entity using random sampling or powerscaled DWS from the corresponding perturbation set -$\mathcal{E}_{\textrm{-}k}$ (constructed with FDA-drugs or CTD-diseases datasets), which is equivalent to applying a type constraint to the replacement entity. For each question to attack, the probability weights in Eq. (3) were calculated for all members of -$\mathcal{E}_{\textrm{-}k}$ using the cosine distance as $h$ between the key and the replacement entities, $k$ and $e_{j}$, in a surrogate word embedding of a BERT model finetuned on biomedical text (Yuan et al., 2022),
$h(k,e_{j})=1-\frac{\text{Embed}(k)\cdot\text{Embed}(e_{j})}{\|\text{Embed}(k)\|\cdot\|\text{Embed}(e_{j})\|}.$ (5)
Using a shared embedding allows model evaluation by the same standard and simplifies the identification of transferable adversarial entities between models, as will be discussed in Section 5.
4.2 Attack execution on LLMs
We selected both generalist (GLMs) and specialist (BLMs) models for benchmarking sampling-based attacks, which require only API access to the LLM. For GLMs, we used instruction-finetuned T5 (Flan-T5) series of models (Chung et al., 2022) and UL2 model (Tay et al., 2022) (Flan-UL2), OpenAI’s GPT-3.5 (text-davinci-003) (Brown et al., 2020) and Google’s PaLM2 (text-bison-001) (Anil et al., 2023). For BLMs, we used MedAlpaca-7B (Han et al., 2023) and Writer’s Palmyra-Med-20B (Kamble and Alshikh, 2023). The selected models contain diverse encoding schemes and architectures. Some of them have appeared on performance leaderboards of related tasks.
We assessed the performance of the LLMs at the baseline level and under sampling-based attacks, all in the zero-shot setting. This resulted in our somewhat lower baseline values than the reported model accuracies that are often achieved with the help of few-shot prompting and prompt optimization to improve performance. All models were evaluated at zero temperature or in the non-sampling setting. The prompts were structured as in Appendix D. Because only up to two model queries are required for each question, the computation is much more affordable than search-based approaches, which can take thousands or more queries (Alzantot et al., 2018). For the attack based on powerscaled DWS, we tuned the hyperparameter $n$ within interval of [-50, 50] using a grid search. The interval was chosen for convenience because beyond this range, the performance drop that the attack incurred on the model doesn’t deviate much from that at or within the interval boundaries.
Drug-mention questions Disease-mention questions
 Model (size)
NoPtb
@@ 4.2 Attack execution on LLMs
 Flan-T5-base (0.25B)
0.276
0.238 (13.8)
0.230 (16.7) / 0.235 (14.9)
0.260
0.226 (13.1)
0.203 (21.9) / 0.221 (15.0)
 Flan-T5-large (0.78B)
0.287
0.252 (12.2)
0.245 (14.6) / 0.243 (15.3)
0.294
0.249 (15.3)
0.230 (21.8) / 0.224 (23.8)
 Flan-T5-xl (3B)
0.318
0.285 (10.4)
0.275 (13.5) / 0.284 (10.7)
0.335
0.298 (11.0)
0.281 (16.1) / 0.293 (12.5)
 Flan-T5-xxl (11B)
0.323
0.292 (9.6)
0.279 (13.6) / 0.287 (11.1)
0.345
0.303 (12.2)
0.276 (20.0) / 0.301 (12.8)
 Flan-UL2 (20B)
0.382
0.349 (8.6)
0.335 (12.3) / 0.342 (10.5)
0.396
0.356 (10.1)
0.334 (15.7) / 0.344 (13.1)
 GPT-3.5 (175B)
0.438
0.398 (9.1)
0.389 (11.2) / 0.391 (10.7)
0.453
0.414 (8.6)
0.401 (11.5) / 0.413 (8.8)
 MedAlpaca-7B
0.460
0.431 (6.3)
0.421 (8.5) / 0.416 (9.6)
0.488
0.439 (10.0)
0.423 (13.3) / 0.436 (10.7)
 Palmyra-Med-20B
0.463
0.413 (10.8)
0.405 (12.5) / 0.414 (10.6)
0.500
0.441 (11.8)
0.425 (15.0) / 0.441 (11.8)
 PaLM2 (350B)
0.287
0.263 (8.4)
0.254 (11.5) / 0.260 (9.4)
0.328
0.302 (7.9)
0.283 (13.7) / 0.297 (11.0)
[…]
Drug-mention questions Disease-mention questions
 Model (size)
NoPtb
@@
 Flan-T5-xl (3B)
0.265
0.227 (14.3)
0.213 (19.6) / 0.225 (15.1)
0.261
0.199 (23.8)
0.194 (25.7) / 0.178 (31.8)
 Flan-T5-xxl (11B)
0.291
0.254 (12.7)
0.234 (19.6) / 0.244 (16.2)
0.376
0.319 (15.2)
0.276 (26.6) / 0.286 (23.9)
 Flan-UL2 (20B)
0.416
0.358 (13.9)
0.353 (15.1) / 0.363 (12.7)
0.408
0.344 (15.7)
0.304 (25.5) / 0.336 (17.6)
 MedAlpaca-7B
0.433
0.383 (11.5)
0.381 (12.0) / 0.381 (12.0)
0.423
0.377 (10.9)
0.357 (15.6) / 0.374 (11.6)
 Palmyra-Med-20B
0.675
0.602 (10.8)
0.585 (13.3) / 0.595 (11.9)
0.661
0.574 (13.2)
0.505 (23.6) / 0.548 (17.1)
Table 2: Adversarial testing of model performance on MedMCQA dataset (same structure as Table 1)
5 Attack analysis and diagnostics
@@
We conducted failure mode analysis using the attack statistics across models and the sampling hyperparameter and through the effects on model explainability. The following summarizes the results.
Two regimes of adversarial entities.
@@
The embedding space picture of the substitution attack in Fig. 1a provides a geometric interpretation of powerscaled DWS to examine different types of adversarial examples by tuning $n$. We use the Gini-Simpson index (Rao, 1982), $\eta_{\text{GS}}=1-\sum_{j}q_{j}^{2}$, to quantify the diversity of adversarial examples and a low value indicates less diversity. Here, $q_{j}\in(0,1]$ is the probability of occurrence of the entity $e_{j}$ among all instances of adversarial entities obtained from sampling, and $\sum_{j}q_{j}=1$. The diversity index indicates the reusability of adversarial entities which we use as a proxy indicator for the universal adversarial attack regime (Moosavi-Dezfooli et al., 2017; Wallace et al., 2019a). By combining the accuracy and the diversity relations as $n$ varies (Fig. 2), we uncover two regimes: (i) the adversarial entities semantically close to ($n\ll 0$) the key $k$ have more diversity; (ii) the adversarial entities semantically far from ($n\gg 0$) the key $k$ are more reusable. These two regimes both lead to an increase in successful attacks. Fig. 2 shows that a larger positive $n$ in Eq. (3) samples more universal adversarial entities than in other situations, leading to a marked drop in the diversity index. We call this behavior the two-regime effect and it is observed in almost all models benchmarked on both QA datasets in varying magnitude.
@@
Figure 3: Model robustness evaluated using R-CAP ($\Delta\text{Acc}/\text{Acc}$) on disease-mention questions from (left) MedQA-USMLE and (right) MedMCQA datasets. The horizontal order partially reflects model performance without perturbation. The colors distinguishes entity-level perturbations by different sampling methods. Perturbations by random sampling (RdPtb) are in purple. Perturbations by powerscaled DWS (DistPtb) in the $n\ll 0$ and $n\gg 0$ regimes are in blue and grey, respectively. A lower value of $\Delta\text{Acc}/\text{Acc}$ indicates a more robust model.
@@
The two-regime effect summarizes the observation that sampling with probability weights either proportional (positive $n$) or inversely proportional (negative $n$) to distance $h$ produces adversarial entities more effectively than random sampling, as shown in Tables 1-2. The benchmarks show that the $n\ll 0$ regime tends to have a more detrimental influence on model performance than the $n\gg 0$ regime. Nevertheless, there exist prominent examples that go against the trend, such as MedAlpaca (for drug-mention questions) and Flan-T5-large in Table 1 and Flan-T5-xl (for disease-mention questions) in Table 2. At small distances, the success of the attack is related to semantic proximity (Mozes et al., 2021). Whereas at larger distances, the attack succeeds potentially due to the obscurity of the sampled entities in the question context, which may produce a profound impact on the model performance (Li et al., 2023). The reduction in the diversity index, $\eta_{\text{GS}}$, shows the dominance of a few replacement entities in successful attacks. We found the most reusable adversarial entities from ranking their occurrences in successful attacks by powerscaled DWS. For drug names, the top-ranked adversarial entities include n-acetylglucosamine (an amino sugar and anti-inflammatory drug)666n-acetylglucosamine on RxList and technetium Tc 99m exametazime (a radiopharmaceutical and contrast agent)777technetium Tc 99m exametazime on RxList. For disease names, some of the chromosome deletion syndromes and rare diseases names are more reusable than others. These entities dominate the sampling results in the $n\gg 0$ regime for all LLMs, indicating their transferability and universality. An example attack using n-acetylglucosamine as recommended medication is in Fig. 4. Adversarial entities in the $n\ll 0$ regime have improper spellings such as missing possessives or they are insufficiently detailed.
Size is not all for model robustness.
A typical metric for model robustness assessment is the change in accuracy before and after perturbation (Zhang et al., 2022). To enable comparison between datasets, we quantify the model robustness through its relative change in accuracy (Acc) upon perturbation (R-CAP), as shown partly in Fig. 3.
$\frac{\Delta\text{Acc}}{\text{Acc}}=\frac{\Delta(n_{\text{TP}}+n_{\text{TN}})}{n_{\text{TP}}+n_{\text{TN}}}$ (6)
R-CAP is equivalently the relative change in the number of correct model predictions (numbers of true positives $n_{\text{TP}}$ and true negatives $n_{\text{TN}}$) (Wang et al., 2023b). Tables 1-2 show that zero-shot adversarial sampling already incurs about 10-20% of relative change in accuracy. For the Flan-T5 series (Chung et al., 2022), R-CAP is more pronounced for smaller and less performant models, indicating the improvement of robustness as models scale up. However, the BLMs are generally more robust to entity substitution than GLMs for models of a similar size. The trend is consistent in both of the two regimes found through hyperparameter tuning of the powerscaled DWS. It is worth noting that although Palmyra-Med-20B (Kamble and Alshikh, 2023) has about three times the number of parameters than MedAlpaca-7B (Han et al., 2023), it is much more sensitive to entity perturbation. Therefore, the smaller BLM is more robust against perturbation. Another observation is that the R-CAP for models upon adversarial perturbations is asymmetrical. For most of the models, sampling adversarial entities closer to the key induces more performance drop than sampling further away.

### v2

@@ 4.1 Dataset selection
To relate to real-world scenarios, we sourced the vocabulary datasets of drug and disease names from existing public databases. The drug names dataset (FDA-drugs) contains over 2.3k unique entities from known drugs approved by the United States Food and Drug Administration (FDA) and curated by Drug Central [^5] https://drugcentral.org/download (Ursu et al., 2017). FDA is a world authority that approves the legal distribution of drugs, therefore the dataset represents the drug-entity names encountered in daily life. The disease names dataset [^6] https://ctdbase.org/downloads/ (CTD-diseases) contains over 9.8k unique entities from the Comparative Toxicogenomics Database (Davis et al., 2009), which contains a comprehensive collection of uniquely documented chemical-gene-disease interactions for humans. Both databases are curated by domain experts through regular updates, and unlike traditional text corpora (Mohan & Li, 2019), the entities in these two datasets contain no redundancy. The 2023 version of the datasets was used after minor data processing.
@@
We selected over 9.3k questions from the MedQA-USMLE (Jin et al., 2021) dataset and over 3.8k questions from the MedMCQA (Pal et al., 2022) dataset for benchmarking. Both datasets are entity-rich and cover a wide range of biomedical specialties and topics. The MedQA-USMLE dataset contains long-context questions that are used to assess medical students in the United States. The MedMCQA dataset contains short questions used in medical exams in India. They have been used in recent works for assessing the biomedical knowledge in LLMs (Singhal et al., 2023a; Han et al., 2023; Singhal et al., 2023b). Both datasets are publicly available and don’t contain personal information. We annotated the NEs in the two datasets according to the entity types of the Unified Medical Language System (UMLS) (Bodenreider, 2004) using scispaCy (Neumann et al., 2019). We then divided the QA datasets into drug- or disease-mention questions by the binary mention labels constructed according to the UMLS entity types (see Appendix C).
4.2 Type-consistent entity substitution
Implementing attacks that target entities requires accounting for type consistency and compatibility with simultaneous multi-word (or span-level) operations, because of the inseparable nature of the entity components (e.g. growth hormone, acquired immunodeficiency syndrome).
Algorithm 1 Type-consistent entity substitution (TCES) template for collecting adversarial entities in distractors.
Inputs: Question Q, LLM, entity type $\tau$, query budget $B$.
Internals: Number of choices $N_{\text{ch}}$, text t, entity Ent, embedding Emb, correct answer k, and evaluator function $\mu$.
Outputs: Number of queries and the replacement entity, None if unsuccessful when the query budget is exhausted.
 TCESAttacker(Q, Model, $\tau$, $B$)
[…]
     for $c\leftarrow 1$ to $N_{\text{ch}}$ do
@@ 4.2 Type-consistent entity substitution
     $i\leftarrow 0$
     while $i<B$ do
@@
[…]
         $i\leftarrow i+1$
     return
This effectively results in more substitution patterns (one to two words, two to one word, two to two words, two to three words, etc.) than the more common one-to-one synonym substitution (Mozes et al., 2021). We describe type-consistent entity substitution (TCES) in Algorithm 1 as the task-agnostic attack template. TCES is a general and contextualized attack for entities, with the goal of adversarially modifying model outputs with only one span-level substitution, implemented using token boundaries. The template can accommodate single or multi-query attacks and it is applicable to any task that may be formatted into QA (Gardner et al., 2019).
[…]
4.3 Attack execution on LLMs
Victim LLMs
We selected both generalist (GLMs) and specialist (BLMs) models for robustness evaluation via blackbox adversarial attacks. The criterion we adopted here for a model to have biomedical knowledge is that it should have a baseline performance better than random guessing (e.g. an accuracy of 0.25 for a specific four-choice QA task) in the relevant domain-specific task. For GLMs, we used instruction-finetuned T5 (Flan-T5) series of models (Chung et al., 2024) and UL2 model (Tay et al., 2022) (Flan-UL2). For BLMs, we used MedAlpaca-7B (Han et al., 2023), Llama2-Medtuned-13B (LM2-Medtuned-13B) (Rohanian et al., 2024) and Writer’s Palmyra-Med-20B (Kamble & Alshikh, 2023). The selected models contain diverse encoding schemes and architectures and their sizes are $\sim$ 1-20B.
Attack settings
To single out the effect of the attacks, all models were evaluated at zero temperature or in the non-sampling setting and model inference was conducted in the zero-shot setting with only basic prompt instructions (see the prompt structure in Appendix D). Model inference of Palmyra-Med-20B (Kamble & Alshikh, 2023) used 4-bit quantization to improve speed (Dettmers & Zettlemoyer, 2023). The evaluation settings resulted in our somewhat lower baseline accuracies than the reported ones which are often achieved with the help of few-shot prompting and prompt optimization. The DiscreteZOO attack (Berger et al., 2021) was originally implemented in the textattack framework (Morris et al., 2020b), which we updated to be compatible with span-level perturbation using its boundaries, multi-GPU split-model inference, and decoder-based billion-parameter LLMs by modifying the attack recipe. To demonstrate the generality of the attack settings, we focus on the following three aspects in the benchmarks.
  • Query budget: We used a fixed query budget for three types of model evaluations: (i) Single-query (budget = 1) sampling-based attacks were used as the reference because the DiscreteZOO attack requires a minimum of 3 model queries; (ii) Multi-query attacks used a budget of 8 for reasonable computational cost across all models and attack settings for both sampling- and search-based attacks; (iii) The query scaling trends of specific LLMs and different attack settings were investigated with fixed budgets under 100 per input instance.
  • Text embedding: Both PDWS and the DiscreteZOO attacks require sampling from a text embedding. The counter-fitted GloVe embedding (Mrkšić et al., 2016) was a default for DiscreteZOO, but GloVe is poor for representing specialized vocabulary in the biomedical domain (Wang et al., 2018). Therefore, here we chose two recent embedding models that offer an improved latent representation of text relations for the benchmarks: (i) CODER (Yuan et al., 2022), a BERT-structured embedding model contrastively pretrained to distinguish UMLS knowledge graph concepts; (ii) GTE-base (Li et al., 2023b) from constrastive learning of general web-scale text data.
  • Hyperparameters: For single-query PDWS attacks, we tuned the primary hyperparameter $n$ within the interval of [-50, 50] using grid search with a step of 5 or 10 because the local maxima of ASR appear on the positive and negative sides. For multi-query PDWS attacks, hyperparameter $n$ was briefly tre-uned around its optimal value in the single-query attack. Most tuned hyperparameters fall within [-30, -5] and [5, 30]. Operating the DiscreteZOO attack in a compatible setting for entities requires two key modifications: (i) Use a random word ranking, which allows the attack algorithm to operate in the low-query-budget setting. (ii) Disable the similarity threshold for candidate point selection in gradient estimation because only entity type consistency is required. The ASR was greatly improved after this step. Benchmarks of the DiscreteZOO attacks were carried out with these settings.
5 Finding and assessing adversarial entities
@@
We conducted failure mode analysis using the attack statistics gathered across models and hyperparameter settings. The following analyses compare methods to identify adversarial entities through query-efficient attacks and uncover their characteristics, including the distribution in the embedding space and their effects on model explainability.
Figure 2: Powerscaled DWS of adversarial distractors exhibits a two-regime effect at negative and positive $n$ values (see Eq. 3) in model accuracy (top) and diversity index (bottom) of replacement entities in successful attacks. The observed similar behaviors are compared across models and datasets in (a) Flan-T5-xxl on MedQA-USMLE, (b) Palmyra-Med-20B on MedQA-USMLE, (c) Flan-T5-xl on MedMCQA, (d) Flan-UL2 on MedMCQA. Disease and drug-mention questions are separated by colors. The vertical dashed line indicates the location of random sampling.
{NiceTabular}
@|l @ P1cm @ |P1.2cm @ P1.2cm @ P2.4cm @ |P1.2cm @ P1.2cm @ P2.4cm |@ MedQA-USMLE MedMCQA
Model (size) Entity
mention Baseline
(Acc) RandS@@@@1
(ASR %) PDWS@@@@1 (-/+)
+CODER (ASR %) Baseline
(Acc) RandS@@@@1
(ASR %) PDWS@@@@1 (-/+)
+CODER (ASR %)
Flan-T5-xl (3B) Drugs 0.318 10.4 13.5 / 10.7 0.265 14.3 19.6 / 15.1
Diseases 0.335 11.0 16.1 / 12.5 0.261 23.8 25.7 / 31.8
Flan-T5-xxl (11B) Drugs 0.323 9.6 13.6 / 11.2 0.291 12.7 19.6 / 16.2
Diseases 0.345 12.2 20.0 / 12.8 0.376 15.2 26.6 / 23.9
Flan-UL2 (20B) Drugs 0.382 8.6 12.3 / 10.5 0.416 13.9 15.1 / 12.7
Diseases 0.396 10.1 15.7 / 13.1 0.408 15.7 25.5 / 17.7
MedAlpaca-7B Drugs 0.474 7.4 9.3 / 10.3 0.433 11.6 13.2 / 12.7
Diseases 0.488 9.9 13.3 / 10.7 0.436 13.5 18.1 / 14.2
LM2-Medtuned-13B Drugs 0.371 16.4 18.1 / 17.8 0.568 15.3 17.3 / 18.1
Diseases 0.398 19.6 23.1 / 19.6 0.515 19.0 27.6 / 21.0
Palmyra-Med-20B Drugs 0.382 12.6 13.3 / 13.5 0.614 10.2 13.8 / 12.7
Diseases 0.441 16.9 21.5 / 18.7 0.624 15.4 20.1 / 11.9
Table 1: Single-query (budget = 1) attack success rates (ASRs) on subsets of MedQA-USMLE and MedMCQA datasets by perturbing the drug or disease entity mentions. Zero-shot baseline (Baseline) accuracy is included for reference next to the ASR of sampling attacks using perturbation by random sampling (RandS), powerscaled distance-weighted sampling (PDWS). The “-” and “+” signs indicate the two regimes ($n<0$ and $n>0$). For each model, the highest ASR for a type of entity-mention question in each dataset is bolded.
5.1 Two regimes of adversarial entities
Diversity quantification of adversarial entities
@@ 5.1 Two regimes of adversarial entities
The embedding space picture of the substitution attack in Fig. 1a provides a geometric interpretation of powerscaled DWS to examine different types of adversarial examples by tuning $n$. We use the Gini-Simpson index (Rao, 1982), $\eta_{\text{GS}}=1-\sum_{j}q_{j}^{2}$, to quantify the diversity of adversarial examples and a low value indicates less diversity. Here, $q_{j}\in(0,1]$ is the probability of occurrence of the entity $e_{j}$ among all instances of adversarial entities obtained from sampling, and $\sum_{j}q_{j}=1$. The diversity index indicates the reusability of adversarial entities which we use as a proxy indicator for the universal adversarial attack regime (Moosavi-Dezfooli et al., 2017; Wallace et al., 2019a).
The dependence of the ASR and diversity on the hyperparameter $n$ in Fig. 2 indicates two regimes: (i) the adversarial entities semantically close to ($n<0$) the key $k$ have a greater diversity; (ii) the adversarial entities semantically far from ($n>0$) the key $k$ are more reusable. TCES in these two regimes both can lead to an increase in successful attacks from random sampling (equivalent to $n=0$), as shown in Tables 5-5.1. Fig. 2 shows that a larger positive $n$ in Eq. 3 samples more universal adversarial entities than in other situations, leading to a marked drop in the diversity index. Moreover, there exist local maxima of ASR in both regimes with finite $n$, which were discovered by hyperparameter tuning. We call these behaviors the two-regime effect and they were observed in almost all models evaluated here in varying magnitudes for both QA datasets, thanks to the controllable spatial coverage of sampling methods incorporating distance information. In both single-query and multi-query attacks, the benchmarks show that the $n<0$ regime of PDWS attacks tends to have a greater influence on model performance than the $n>0$ regime, although prominent counterexamples do exist (MedAlpaca-7B and Llama2-Medtuned-13B (for drug-mention questions), and Flan-T5-xl (for disease-mention questions) in Table 5).
@@
Figure 3: Model robustness from single-query adversarial attacks using results in Table 5. The models (GLMs and BLMs) were evaluated on drug-mention questions from MedQA-USMLE (left) and disease-mention questions from MedMCQA (right) datasets. The GLMs and BLMs are ordered horizontally by their sizes. The bar colors distinguish between different attack methods. Perturbations by random sampling (RS) are in purple. Perturbations by powerscaled DWS (PDWS) in the $n<0$ and $n>0$ regimes are in blue and grey, respectively.
Adversarial entity characteristics
@@
The two-regime effect indicates that the success of the attack may come from two sources. At small distances, semantic proximity (Mozes et al., 2021) reduces the distinguishability between the key and the distractor, therefore increasing the chance of model failure. Whereas at larger distances, the attack succeeds potentially due to the obscurity of the distractor entities in the question context, which can also profoundly impact the model performance (Li et al., 2023a). This explanation may be further supported by examining the diversity index, $\eta_{\text{GS}}$, which shows the dominance of a few replacement entities in successful attacks at large distances. We found these highly reusable adversarial entities by ranking their occurrences in successful attacks by powerscaled DWS. For drug names, the top-ranked adversarial entities include n-acetylglucosamine (an amino sugar and anti-inflammatory drug)777n-acetylglucosamine on RxList and technetium Tc 99m exametazime (a radiopharmaceutical and contrast agent)888technetium Tc 99m exametazime on RxList. For disease names, some of the chromosome deletion syndromes and rare diseases names are more reusable than others. These entities dominate the sampling results in the $n>0$ regime for all LLMs, indicating their transferability and universality.
{NiceTabular}
@|l @ P1cm @ |P1cm @ P1.8cm @ P1.3cm @ P1.3cm @ |P1cm @ P1.8cm @ P1.3cm @ P1.3cm |@ MedQA-USMLE (ASR %) MedMCQA (ASR %)
Model (size) Entity
mention RandS
@@@@8 PDWS@@@@8 (-/+)
+CODER DZOO@@@@8
+CODER DZOO@@@@8
+GTE-base RandS
@@@@8 PDWS@@@@8 (-/+)
+CODER DZOO@@@@8
+CODER DZOO@@@@8
+GTE-base
Flan-T5-xl (3B) Drugs 22.0 29.3 / 23.3 21.7 19.5 32.5 48.3 / 36.2 34.3 28.2
Diseases 27.2 38.8 / 24.8 20.9 23.2 47.1 61.7 / 49.0 45.8 39.3
Flan-UL2 (20B) Drugs 22.5 27.5 / 22.3 20.4 19.2 26.4 35.6 / 22.4 30.7 26.4
Diseases 25.8 34.9 / 26.8 25.7 24.0 39.2 49.5 / 31.6 41.3 37.0
MedAlpaca-7B Drugs 20.3 22.8 / 24.5 22.8 20.5 24.9 34.2 / 27.5 42.6 36.8
Diseases 25.2 30.1 / 21.6 26.5 25.2 31.2 41.7 / 25.0 40.5 38.1
Palmyra-Med-20B Drugs 30.1 31.4 / 31.3 29.1 29.6 28.5 37.3 / 31.9 35.4 35.0
Diseases 33.2 35.1 / 30.2 32.3 34.5 41.1 48.4 / 39.6 46.3 50.1
Table 2: Multi-query (here budget = 8) attack success rates (ASR %) on various LLMs by perturbing the drug and disease entity mentions in MedQA-USMLE and MedMCQA datasets. The PDWS attack uses the embedding from CODER and DiscreteZOO (DZOO) attacks use either CODER or GTE-base to calculate the cosine distance. For each model, the highest ASR for a type of entity-mention question in each dataset is bolded.
5.2 Scaling characteristics from adversarial entities
Size is not all for scaling model robustness.
We compare model robustness using single-query ASR as the metric in Table 5 and Fig. 3. For the Flan-T5 series of GLMs (Chung et al., 2024) and Flan-UL2 (Tay et al., 2022), the single-query ASR is more pronounced in smaller and less performant models, indicating the improvement of robustness as the model’s size grows. However, an opposite trend of robustness is seem in BLMs. It is worth noting that although Palmyra-Med-20B (Kamble & Alshikh, 2023) has about three times the number of parameters than MedAlpaca-7B (Han et al., 2023), the larger model is much more sensitive to entity perturbation. In terms of ASR, the BLMs are comparable to the GLMs in each type of questions evaluated on. Another observation is that the ASR for models is asymmetrical under PDWS attacks. For most of the models, sampling adversarial entities with inverse-distance-proportional weights ($n<0$ in Eq. 3) induces more performance drop than sampling with distance-proportional weights ($n>0$ in Eq. 3), yet prominent counterexamples exist. The attack performance is further investigated in the sequential, multi-query setting (budget = 8) in Table 5.1, where we also compared sampling- against search-based DiscreteZOO attacks. Here, PDWS attacks in the $n<0$ regime are still largely the most effective.
Figure 4: Example scaling curves of the query budget against ASR for disease-mention questions with (a) Flan-T5-xl on MedMCQA dataset and (b) MedAlpaca model on MedQA-USMLE dataset. The curves are generated using TCES of disease names with the methods described in the legends. The DiscreteZOO attacks were run in the low-query-budget setting. Executing the random sampling (RS) attack doesn’t require a text embedding, while the other attacks were evaluated with the CODER or GTE-base as text embedding.
Sampling in specialized text embedding has favorable query scaling over blackbox gradient attacks.
A query scaling study of adversarial attacks investigates attack success with varying query budgets per input instance (Shukla et al., 2021). It is a comprehensive evaluation of attack methods, but not yet frequently adopted in most publications, partly due to the computational cost. Fig. 4 compares the query scaling of both sampling- and search-based methods for TCES. For low-query-budget attacks (Fig. 1a), each scaling curve exhibits a rapid rise phase and a plateau phase, separated by an inflection point. The key message here is that the advantage of PDWS attacks over DiscreteZOO exists with a choice of text embedding. It is greater in specialized text embeddings but is reduced or diminished in general ones thanks to the difference in neighborhood structures. We expand the analysis in the following in three directions.
  • Query scaling depends on text embedding: Changing the text embedding affects the query scaling behavior of both PDWS and DiscreteZOO attacks. Fig. 4 shows that the influence of the embedding choice on PDWS is more pronounced in the rapid rise phase, while it influences the DiscreteZOO primarily in the plateau phase. Specifically, changing the text embedding from domain-specialized CODER (Yuan et al., 2022) to the general GTE-base (Li et al., 2023b), the performance advantage of PDWS over DiscreteZOO attacks shrunk in Flan-T5-xl (Fig. 4a) and inverted in MedAlpaca-7B (Fig. 4b).
  • Sampling hyperparameter controls scaling curve shape in PDWS attacks: The hyperparameter $n$ changes the inflection point of the scaling curve. When the query budget is past the inflection point of the curve, random sampling starts to surpass PDWS ($n>0$) in ASR. This is a consequence of the distribution of adversarial entities in the embedding space and the regional bias derived from the sampling weights. The examples in Fig. 4 show that the largest performance gap between the PDWS and DiscreteZOO attacks appears at around the inflection point of the respective query scaling curve.
  • Gradient quality affects query scaling: The query complexity in blackbox gradient estimation is a long-standing theoretical topic in derivative-free optimization (L’Ecuyer & Yin, 1998) and online learning (Duchi et al., 2015). Nevertheless, in application settings such as adversarial machine learning, how gradient quality affects query scaling remains less studied for text data. Gradient estimation consumes a part of the total query budget, leading to a tradeoff between the gradient quality and terminal ASR attainable in the query scaling curve. We empirically explored the query scaling behavior of the DiscreteZOO attack in Appendix E with different candidate points in the multi-point gradient estimation, which in effect alters the gradient quality. The results show that having too few candidate points (i.e. small $M$ in Eq. 6) leads to an early onset of the plateau phase, thereby limiting the highest attainable ASR within the total query budget, whereas having many candidate points can lead to a slower growth of ASR in the rapid rise phase of the query scaling curve.